OpenStudy (anonymous):

How do you determine the end behavior of polynomial functions without a graph?

OpenStudy (anonymous):

For example: \[y=-6x ^{3}+9x ^{2}+4x-7\]

OpenStudy (anonymous):

the degree is odd, and the leading coefficient is negative

OpenStudy (anonymous):

that means as \(x\to -\infty\) we know \(y\to \infty\) and as \(x\to \infty\) then \(y\to -\infty\)

OpenStudy (anonymous):



OpenStudy (anonymous):

okay here is a better picture http://www.wolframalpha.com/input/?i=-6x+^ {3}%2B9x+^{2}%2B4x-7